OpenStudy (anonymous):

Is the sequence –9.9, –13, –16, –19.2, … arithmetic? If so, identify the common difference.

OpenStudy (anonymous):

???

OpenStudy (katrinakaif):

Can you give it a try first?

OpenStudy (anonymous):

-9.9 + ? = -13 -13 + ? = -16 -16 + ? = -19.2 Does the same number replace the "?" in all three? If it does then it's arithmetic and that number is the common difference. If not, it's not arithmetic.